# MOSAIC: Modeling Social AI for Content Dissemination and Regulation in Multi-Agent Simulations

### v2

@@ A.4 Do Content Topics Matter?
Following topic assignment, we conducted a detailed statistical analysis. Engagement metrics—including mean, median, and standard deviation for likes, shares, comments, and overall engagement scores—were calculated for each topic. To statistically assess whether variations in engagement across topics were significant, we performed an ANOVA (Analysis of Variance).
@@ A.5 Clues from Agents’ Own Reasoning Traces and Recommender System
The lack of clear correlation between user profiles, content topics, or temporal properties, and engagement patterns suggested that maybe the way we present feed to the agents has an influence on what content ends up being popular. Here we discuss our feed prioritization algorithm. Our simulation does not employ a sophisticated recommender system. Our feed prioritization in the simulation relies primarily on recency and existing follow relationships, rather than explicit engagement metrics such as likes or shares. Posts, regardless of whether they’re from followed or non-followed users, are generally ordered based on creation time, ensuring that newer content receives greater visibility. However, content from followed users gains additional prioritized exposure due to dedicated allocations in the feed. This structure might create a follow-based feedback loop: when User B follows User A, A’s posts consistently appear in B’s feed, enhancing A’s opportunities for engagement through likes, comments, and shares. Higher engagement subsequently boosts A’s visibility to other users who view these interactions, increasing the likelihood of additional follows and further amplifying this cycle.
@@
Vocabulary analysis further emphasizes these distinctions, revealing specialized linguistic patterns for each type of engagement. Flagging content uses specific moderation-related language such as "misinformation," "harmful," and "credible," whereas community-oriented engagements like sharing, liking, and commenting frequently reference concepts like "community," "support," and "alignment." Following actions highlight terms related to content curation and long-term value, including "consistently," "valuable," and "insights."
@@ Appendix B Persona Generation Details
Here we describe the questions that we sampled and generated for the agent users. The generated personas are stored in JSONL format, with each entry containing a unique identifier, a descriptive narrative, and associated behavioral labels.
@@ B.1 Persona Replication from Human Survey
The persona generation method begins by transforming structured survey responses collected from Prolific participants into rich, natural language character descriptions suitable for use in agent-based simulations. Each participant’s responses — covering a wide range of personal, demographic, social, and psychological traits—are encoded in JSONL format, where each line corresponds to a different individual. We first this file into a list of Python dictionaries, each representing a single participant’s answers. The preprocessing pipeline then embeds each answer into a templated sentence structure. This includes details such as age, gender, residential background, number of places lived, favorite activities, values, political stance, income, ethnicity, language, education, religion, social tendencies, hobbies, relationship values, personality, future goals, significant life events, friendship values, and hypothetical financial decisions. By expressing these traits in fluent, first-person-style English, the function essentially replicates each participant’s worldview and identity into a lifelike persona that can guide agent behavior in social simulations. In the final step iterates through all participant entries, generates the corresponding natural language persona for each one, and writes the enriched data—including both the original responses and the generated description—back into a new JSONL file. This process creates a bridge between raw human survey data and psychologically grounded agent profiles, enabling more realistic and diverse behaviors in multi-agent environments.
@@ B.3 User Generation and Instantiation
As illustrated in Fig. 1, personas are generated using a combination of predefined questions and sampling from probabilistic distributions stored in an agent_bank profile collection, inspired by Park et al. (2024). Key demographic attributes such as age, gender, ethnicity, and primary language are assigned probabilistically to mirror real-world distributions. For instance, age follows a normal distribution centered around 35 years, while other attributes are sampled based on predefined probabilities. We disclose these questions and describe more detail of the methodology in Appendix B. After synthesizing the structured profiles for agents, we construct a natural language description for each of them. This process leverages a mixture of deterministic rules and LLM-based augmentation using GPT-4o Hurst et al. (2024) to enhance diversity and realism.
@@
Once personas are generated, they are instantiated as agent users within the simulation. Each agent is assigned a unique user ID and a persona profile that includes background details and interest labels. The relational database serves as the backbone for recording agent activities, ensuring persistent storage of interactions, post engagements, and behavioral updates. This database facilitates dynamic user tracking and enables post-simulation analysis of engagement trends and content spread. We provide more implementation details in Appendix F.
@@ Appendix D Details of the Human Study and IRB
The study was open to 20,240 eligible participants from a larger Prolific population of 232,330, and we collected 204 valid responses from eligible participants. The survey was conducted via Prolific to collect responses from U.S.-based participants fluent in English. Participants were asked to complete a 12-minute survey assessing their demographic characteristics and social media interactions. The survey, hosted on Google Forms, required no software downloads or special device features and was accessible via mobile, tablet, or desktop. Participant recruitment applied custom screening for language, political spectrum, vaccine opinion, and prior participation, ensuring a targeted sample. Responses were collected using Prolific ID via a question at the start of the form, and participants received a completion code upon finishing. Compensation was set at \$2.40 per participant, equivalent to \$12.00/hour, and submissions were manually reviewed before approval. The median completion time was approximately 14.5 minutes. All members on our research team have obtained IRB approval before the human study was conducted. Our study costs a total of \$480 for participant payment and \$160 of platform fee. Fig. 6 shows the complete breakdown of the 9 key demographic distributions of the 204 human participants. This study was performed under approval from the appropriate institutional ethics review board. Full IRB documentation will be made available upon request.
@@ Appendix E An Extended Version of Related Work
Behavioral Economics and Persuasian Games.
@@
The emergence of large language models (LLMs) has significantly advanced the capabilities of agent-based social simulations, enabling more sophisticated, context-aware interactions. Traditional agent-based modeling relied on predefined rule sets and heuristics, limiting adaptability and realism. Early computational social simulations, such as Schelling’s segregation model Schelling (1971), Sugarscape Epstein and Axtell (1996) and NetLogo-based models Wilensky (1999), provided insights into social dynamics but lacked the ability to generate nuanced, context-dependent behaviors.
@@
Third-party fact-checking, typically conducted by organizations such as Snopes [^11] https://www.snopes.com/, PolitiFact [^12] https://www.politifact.com/, or Google’s partnerships with external organizations Raghunath and Malik (2024); Patel (2024), provides authoritative assessments but faces challenges in scalability and timeliness Zannettou et al. (2019); Uscinski and Butler (2013); Marietta et al. (2015). Crowdsourced fact-checking such as X’s Community Notes, [^13] https://communitynotes.x.com/guide/en/ on the other hand, leverages collective intelligence Panizza et al. (2023) but introduces risks related to expertise and susceptibility to group biases Saeed et al. (2022); Pennycook et al. (2021). There is no consensus on which fact-checking approach is more effective, nor is it well-understood how different moderation strategies interact. This study addresses this gap by leveraging LLM-driven simulations to evaluate different fact-checking mechanisms within controlled environments. By testing various moderation strategies in a scalable, repeatable manner, this work provides insights into the comparative efficacy of community-based, third-party, and hybrid fact-checking interventions in mitigating misinformation.
@@ Appendix G Agent Action Space
In our simulated social media environment, each agent—powered by a large language model—is instantiated with a predefined action space that governs its interactions within the platform. These agents are configured with unique personas via a Config (persona) generation module and endowed with a Memory / reflection module that allows them to recall and adapt based on past experiences. The Action space outlines the full spectrum of behaviors an agent can exhibit: they can share posts, comment or like comments, create or like posts, and flag inappropriate content. Additionally, agents can retrieve their feeds, follow or unfollow other users, or ignore content or interactions altogether. These discrete actions simulate realistic user behavior and social dynamics, enabling rich, emergent interactions in the environment.
@@ G.1 Agent Decision-Making Process
Agents make decisions based on a combination of persona-driven heuristics, memory retrival, and reasoning prompts. The AgentPrompts module formulates structured decision prompts, guiding agents through content engagement options such as liking, sharing, or flagging a post. When engaging, agents provide reasoning for their actions, influenced by (1) Personal Beliefs and Persona Traits: Agents weigh content credibility based on their ideological stance and historical preferences, (2) Engagement Signals: Highly engaged posts are more likely to be reshared due to social validation effects, and (3) Fact-Checking Feedback: Agents integrate fact-checking signals into their reasoning, adjusting their trust in flagged content accordingly.
@@ Appendix H Detailed Experiment Configuration
We describe our experimental settings and configurable variables in more detail in this section. In our simulation, we model a dynamic social network of an arbitrary number of (practically in our experiments up to over 200) LLM-driven agents interacting over the course of a number of discrete time steps. The simulation loop follows a structured core cycle that includes initializing the environment, assigning new users probabilistically (though this one could be disabled in certain runs), content creation, feed-based reactions, and periodic reflective updates. Each agent is instantiated from detailed persona descriptions provided via an external JSONL file, and operates using the GPT-4o engine with a decoding temperature of 1.0 to promote diversity in generated responses. Agents can be configured to create original posts independently, or they can be prompted to only respond to posts depending on the setting. Once the simulation environment is initiated, an agent’s feed consists of a mixture of up to a default of 15 posts from followed users and 10 from non-followed users, drawn from a pool that includes up to 20 injected news items per run. Initial social ties are sparse, with a 10% probability of following another user at initialization, and new user addition and follow behaviors are disabled during the simulation. All of the above numbers are configurable. The experiment evaluates one of the four fact-checking intervention modes described in Section 3, combining both third-party and community-based mechanisms. For each step, if a fact-checking agent is enabled, then a number of posts are selected for potential moderation, with fact-checking outputs generated using a low-temperature (0.3) setting and required to include reasoning. Thresholds are specified for flagging and notetaking behavior if moderation is set active. Periodically, agents reflect on their recent interactions, update memory states, and check their internal objectives, offering a framework for studying emergent behavior, information diffusion, and intervention efficacy in artificial societies.
@@ Appendix I Details of the Memory Module
[…]
@@
Figure 8: Top 50 users with highest engagement.
@@
"None", "Community Based": There is little to no interaction with misinformative content. Interaction with factual posts decreases over time.
"Third Party": Initially, there is higher interaction with misinformative content than factual content. This reverses as interaction with misinformation decreases and with factual content increases.
"Hybrid": There is little to no interaction with misinformative content. There is steady increase in interaction with factual content.
@@
"None", "Community Based", "Hybrid": There is little interaction with misinfomrative content. Interaction with factual content starts off strongly and increases over time.
"Third Party": Initially, there is much higher interaction with misinformative content than factual content. Over time, interaction with misinformation decreases and with factual content increases, but interaction with misinformation is still close to factual content interaction.
@@
"None", "Community Based": There is little interaction with both factual and misinformative content.
"Third Party", "Hybrid": There is some interaction with misinformative content, though less than the consistent and increasing interaction with factual content.
@@
"None": All models have very little interaction with misinformative content. DeepSeek is the only model that increases interaction with factual content, while the others decrease over time.
"Community Based": All models continue to have very little interaction with misinformative content. Again, DeepSeek is the only model that increases interaction with factual content over time.
"Third Party": All experiments intially start with a higher level of interaction with misinformation than interaction with factual information. This trend reverses across all models over time. However, the level of difference in interaction with misinformation and factual information is most pronounced in Claude. Other models have less differentiation.
"Hybrid": Claude has the lowest levels of interaction with misinformative content as well as the highest and most consistent levels of interaction with factual content. DeepSeek initially displays the same trend, while experiencing increasing misinformation engagment in later time steps. GPT-4o has constant interaction with different content types, though slightly higher with factual content.
@@
GPT-4o: GPT shows low overall engagement with both factual and misinformative content under the "None" and "Community Based" regimes. This could reflect a more cautious content sharing dynamic when no clear verification cues are available.
@@
Fig. 14 shows Average Comments across all models. In "Hybrid", comments on factual content steadily increase over time, while misinformation receives minimal and flat engagement. In "Community Based" and "None", misinformation remains low and nearly flat, while factual content is slightly higher. In "Third-Party", there is a steady increase in comments on factual content, while comments on misinformative content start high and decrease. Strong moderation (like hybrid) seems to promote public discussion of factual content while limiting commentary on misinformation.
@@
Fig. 10 shows Average Likes across all models. Likes across all fact checking modes have high engagement over time for factual information, with minimal engagement with misinformation. For "Hybrid" and "Third Party" approaches, likes trend up, while for "Community-Based" and "No Fact Checking", likes trend down over time. Interestingly, while there is almost no engagement in other modes, there remains some likes for misinformation in "Third-Party". Hybrid moderation seems to foster the strongest positive sentiment toward factual content with lowest disengagement with misinformative content.
@@
Fig. 11 shows Average Note Rating count across all models. There are a higher amount of note ratings in "Hybrid" than "Community-Based". Additionally, expected behavior is that there are few notes to rate for factual information.
@@
Fig. 12 shows Average Community Note count across all models. In "Hybrid", there is a steady growth of community notes for misinformation the majority of the time steps. In "Community-Based", there is actually a steady decline. There are also no notes for factual information in "Community-Based", while there are few but steadily growing notes in "Hybrid".
Note Ratings and Community Notes suggests that a "Hybrid" approach best stimulates robust feedback loops for users, including interaction with all types of information and interaction with other users’ notes.
@@
Fig. 13 shows the final action possible for agents, Average Share count across all models. The number of average shares is generally low, being less than 1 per post. In this proportion, for "Hybrid", "Community-Based", and "None" approaches there is a fast increase to decrease and constant low in sharing of misinformation. "Hybrid" has gradually increasing shares for factual information, while "Community-Based" and "None" slowly decrease, though "None" has more overall shares. "Third-Party" experiences the most overall shares, and the sharpest increase in shares of factual information over time as well. Overall, strong fact-checking encourages users to amplify factual content, whereas misinformation rarely achieves viral spread across any setting.
Appendix L Prompts
L.1 Post Creation Prompt
@@
L.2 Feed Reaction Prompt

### v3

@@ A.4 Do Content Topics Matter?
Following the topic assignment, we conducted a detailed statistical analysis. Engagement metrics—including mean, median, and standard deviation for likes, shares, comments, and overall engagement scores—were calculated for each topic. To statistically assess whether variations in engagement across topics were significant, we performed an ANOVA (Analysis of Variance).
@@ A.5 Clues from Agents’ Own Reasoning Traces and Recommender System
The lack of clear correlation between user profiles, content topics, or temporal properties, and engagement patterns suggested that maybe the way we present the feed to the agents influences what content ends up being popular. Here, we discuss our feed prioritization algorithm. Our simulation does not employ a sophisticated recommender system. Our feed prioritization in the simulation relies primarily on recency and existing follow relationships, rather than explicit engagement metrics such as likes or shares. Posts, regardless of whether they’re from followed or non-followed users, are generally ordered based on creation time, ensuring that newer content receives greater visibility. However, content from followed users gains additional prioritized exposure due to dedicated allocations in the feed. This structure might create a follower-based feedback loop: when User B follows User A, A’s posts consistently appear in B’s feed, enhancing A’s opportunities for engagement through likes, comments, and shares. Higher engagement subsequently boosts A’s visibility to other users who view these interactions, increasing the likelihood of additional follows and further amplifying this cycle.
@@
Vocabulary analysis further emphasizes these distinctions, revealing specialized linguistic patterns for each type of engagement. Flagging content uses specific moderation-related language such as "misinformation," "harmful," and "credible," whereas community-oriented engagements like sharing, liking, and commenting frequently reference concepts like "community," "support," and "alignment." The "follow-user" action highlights terms related to content curation and long-term value, including "consistently," "valuable," and "insights."
@@ Appendix B Persona Generation Details
Here, we describe the questions that we sampled and generated for the agent users. The generated personas are stored in JSONL format, with each entry containing a unique identifier, a descriptive narrative, and associated behavioral labels.
@@ B.1 Persona Replication from Human Survey
The persona generation method begins by transforming structured survey responses collected from Prolific participants into rich, natural language character descriptions suitable for use in agent-based simulations. Each participant’s responses — covering a wide range of personal, demographic, social, and psychological traits—are encoded in JSONL format, where each line corresponds to a different individual. We first put this file into a list of Python dictionaries, each representing a single participant’s answers. The preprocessing pipeline then embeds each answer into a templated sentence structure. This includes details such as age, gender, residential background, number of places lived, favorite activities, values, political stance, income, ethnicity, language, education, religion, social tendencies, hobbies, relationship values, personality, future goals, significant life events, friendship values, and hypothetical financial decisions. By expressing these traits in fluent, first-person-style English, the function essentially replicates each participant’s worldview and identity into a lifelike persona that can guide agent behavior in social simulations. In the final step, it iterates through all participant entries, generates the corresponding natural language persona for each one, and writes the enriched data—including both the original responses and the generated description—back into a new JSONL file. This process creates a bridge between raw human survey data and psychologically grounded agent profiles, enabling more realistic and diverse behaviors in multi-agent environments.
@@ B.3 User Generation and Instantiation
As illustrated in Fig. 1, personas are generated using a combination of predefined questions and sampling from probabilistic distributions stored in an agent_bank profile collection, inspired by Park et al. (2024). Key demographic attributes such as age, gender, ethnicity, and primary language are assigned probabilistically to mirror real-world distributions. For instance, age follows a normal distribution centered around 35 years, while other attributes are sampled based on predefined probabilities. We disclose these questions and describe more details of the methodology in Appendix B. After synthesizing the structured profiles for agents, we construct a natural language description for each of them. This process leverages a mixture of deterministic rules and LLM-based augmentation using GPT-4o Hurst et al. (2024) to enhance diversity and realism.
@@
Once personas are generated, they are instantiated as agent users within the simulation. Each agent is assigned a unique user ID and a persona profile that includes background details and interest labels. The relational database serves as the backbone for recording agent activities, ensuring persistent storage of interactions, post-engagements, and behavioral updates. This database facilitates dynamic user tracking and enables post-simulation analysis of engagement trends and content spread. We provide more implementation details in Appendix F.
@@ Appendix D Details of the Human Study and IRB
The study was open to 20,240 eligible participants from a larger Prolific population of 232,330, and we collected 204 valid responses from eligible participants. The survey was conducted via Prolific to collect responses from U.S.-based participants fluent in English. Participants were asked to complete a 12-minute survey assessing their demographic characteristics and social media interactions. The survey, hosted on Google Forms, required no software downloads or special device features and was accessible via mobile, tablet, or desktop. Participant recruitment applied custom screening for language, political spectrum, vaccine opinion, and prior participation, ensuring a targeted sample. Responses were collected using Prolific ID via a question at the start of the form, and participants received a completion code upon finishing. Compensation was set at \$2.40 per participant, equivalent to \$12.00/hour, and submissions were manually reviewed before approval. The median completion time was approximately 14.5 minutes. All members on our research team have obtained IRB approval before the human study was conducted. Our study costs a total of \$480 for participant payment and \$160 of the platform fee. Fig. 6 shows the complete breakdown of the 9 key demographic distributions of the 204 human participants. This study was performed under approval from the appropriate institutional ethics review board. Full IRB documentation will be made available upon request.
@@ Appendix E An Extended Version of Related Work
Behavioral Economics and Persuasion Games.
@@
The emergence of large language models (LLMs) has significantly advanced the capabilities of agent-based social simulations, enabling more sophisticated, context-aware interactions. Traditional agent-based modeling relied on predefined rule sets and heuristics, limiting adaptability and realism. Early computational social simulations, such as Schelling’s segregation model Schelling (1971), Sugarscape Epstein and Axtell (1996), and NetLogo-based models Wilensky (1999), provided insights into social dynamics but lacked the ability to generate nuanced, context-dependent behaviors.
@@
Third-party fact-checking, typically conducted by organizations such as Snopes [^11] https://www.snopes.com/, PolitiFact [^12] https://www.politifact.com/, or Google’s partnerships with external organizations Raghunath and Malik (2024); Patel (2024), provides authoritative assessments but faces challenges in scalability and timeliness Zannettou et al. (2019); Uscinski and Butler (2013); Marietta et al. (2015). Crowdsourced fact-checking, such as X’s Community Notes, [^13] https://communitynotes.x.com/guide/en/ on the other hand, leverages collective intelligence Panizza et al. (2023) but introduces risks related to expertise and susceptibility to group biases Saeed et al. (2022); Pennycook et al. (2021). There is no consensus on which fact-checking approach is more effective, nor is it well-understood how different moderation strategies interact. This study addresses this gap by leveraging LLM-driven simulations to evaluate different fact-checking mechanisms within controlled environments. By testing various moderation strategies in a scalable, repeatable manner, this work provides insights into the comparative efficacy of community-based, third-party, and hybrid fact-checking interventions in mitigating misinformation.
@@ Appendix G Agent Action Space
In our simulated social media environment, each agent—powered by a large language model—is instantiated with a predefined action space that governs its interactions within the platform. These agents are configured with unique personas via a configuration (persona) generation module and endowed with a memory/reflection module that allows them to recall and adapt based on past experiences. The Action space outlines the full spectrum of behaviors an agent can exhibit: they can share posts, comment or like comments, create or like posts, and flag inappropriate content. Additionally, agents can retrieve their feeds, follow or unfollow other users, or ignore content or interactions altogether. These discrete actions simulate realistic user behavior and social dynamics, enabling rich, emergent interactions in the environment.
@@ G.1 Agent Decision-Making Process
Agents make decisions based on a combination of persona-driven heuristics, memory retrieval, and reasoning prompts. The AgentPrompts module formulates structured decision prompts, guiding agents through content engagement options such as liking, sharing, or flagging a post. When engaging, agents provide reasoning for their actions, influenced by (1) Personal Beliefs and Persona Traits: Agents weigh content credibility based on their ideological stance and historical preferences, (2) Engagement Signals: Highly engaged posts are more likely to be reshared due to social validation effects, and (3) Fact-Checking Feedback: Agents integrate fact-checking signals into their reasoning, adjusting their trust in flagged content accordingly.
@@ Appendix H Detailed Experiment Configuration
We describe our experimental settings and configurable variables in more detail in this section. In our simulation, we model a dynamic social network of an arbitrary number of (practically in our experiments, up to over 200) LLM-driven agents interacting over the course of a number of discrete time steps. The simulation loop follows a structured core cycle that includes initializing the environment, assigning new users probabilistically (though this one could be disabled in certain runs), content creation, feed-based reactions, and periodic reflective updates. Each agent is instantiated from detailed persona descriptions provided via an external JSONL file, and operates using the GPT-4o engine with a decoding temperature of 1.0 to promote diversity in generated responses. Agents can be configured to create original posts independently, or they can be prompted to only respond to posts, depending on the setting. Once the simulation environment is initiated, an agent’s feed consists of a mixture of up to a default of 15 posts from followed users and 10 from non-followed users, drawn from a pool that includes up to 20 injected news items per run. Initial social ties are sparse, with a 10% probability of following another user at initialization, and new user addition and follow behaviors are disabled during the simulation. All of the above numbers are configurable. The experiment evaluates one of the four fact-checking intervention modes described in Section 3, combining both third-party and community-based mechanisms. For each step, if a fact-checking agent is enabled, then a number of posts are selected for potential moderation, with fact-checking outputs generated using a low-temperature (0.3) setting and required to include reasoning. Thresholds are specified for flagging and note-taking behavior if moderation is set to active. Periodically, agents reflect on their recent interactions, update memory states, and check their internal objectives, offering a framework for studying emergent behavior, information diffusion, and intervention efficacy in artificial societies.
@@ Appendix I Details of the Memory Module
$\text{Relevance}=\text{Importance}\times\text{Decay}$
@@
Figure 8: Top 50 users with the highest engagement.
@@
"None", "Community Based": There is little to no interaction with false news. Interaction with factual posts decreases over time.
"Third Party": Initially, there is higher interaction with false news than factual content. This reverses as interaction with misinformation decreases and with factual content increases.
"Hybrid": There is little to no interaction with false news. There is a steady increase in interaction with factual content.
@@
"None", "Community Based", "Hybrid": There is little interaction with misinformation. Interaction with factual content starts off strongly and increases over time.
"Third Party": Initially, there is much higher interaction with false news than factual content. Over time, interaction with misinformation decreases and with factual content increases, but interaction with misinformation is still close to factual content interaction.
@@
"None", "Community Based": There is little interaction with both factual and false news.
"Third Party", "Hybrid": There is some interaction with false news, though less than the consistent and increasing interaction with factual content.
@@
"None": All models have very little interaction with false news. DeepSeek is the only model that increases interaction with factual content, while the others decrease over time.
"Community Based": All models continue to have very little interaction with false news. Again, DeepSeek is the only model that increases interaction with factual content over time.
"Third Party": All experiments initially start with a higher level of interaction with misinformation than interaction with factual information. This trend reverses across all models over time. However, the level of difference in interaction with misinformation and factual information is most pronounced in Claude. Other models have less differentiation.
"Hybrid": Claude has the lowest levels of interaction with false news as well as the highest and most consistent levels of interaction with factual content. DeepSeek initially displays the same trend, while experiencing increasing misinformation engagement in later time steps. GPT-4o has constant interaction with different content types, though slightly higher with factual content.
@@
GPT-4o: GPT shows low overall engagement with both factual and false news under the "None" and "Community Based" regimes. This could reflect a more cautious content-sharing dynamic when no clear verification cues are available.
@@
Fig. 14 shows Average Comments across all models. In "Hybrid", comments on factual content steadily increase over time, while misinformation receives minimal and flat engagement. In "Community Based" and "None", misinformation remains low and nearly flat, while factual content is slightly higher. In "Third-Party", there is a steady increase in comments on factual content, while comments on false news start high and decrease. Strong moderation (like hybrid) seems to promote public discussion of factual content while limiting commentary on misinformation.
@@
Fig. 10 shows Average Likes across all models. Likes across all fact-checking modes have high engagement over time for factual information, with minimal engagement with misinformation. For "Hybrid" and "Third Party" approaches, likes trend up, while for "Community-Based" and "No Fact Checking", likes trend down over time. Interestingly, while there is almost no engagement in other modes, there remains some likes for misinformation in "Third-Party". Hybrid moderation seems to foster the strongest positive sentiment toward factual content, with the lowest disengagement with false news.
@@
Fig. 11 shows the average Note Rating count across all models. There is a higher amount of note ratings in "Hybrid" than "Community-Based". Additionally, expected behavior is that there are few notes to rate for factual information.
@@
Fig. 12 shows the average Community Note count across all models. In "Hybrid", there is a steady growth of community notes for misinformation for the majority of the time steps. In "Community-Based", there is actually a steady decline. There are also no notes for factual information in "Community-Based", while there are a few but steadily growing notes in "Hybrid".
Note Ratings and Community Notes suggest that a "Hybrid" approach best stimulates robust feedback loops for users, including interaction with all types of information and interaction with other users’ notes.
@@
Fig. 13 shows the final action possible for agents, Average Share count across all models. The number of average shares is generally low, being less than 1 per post. In this proportion, for "Hybrid", "Community-Based", and "None" approaches, there is a fast increase to decrease and constant low in the sharing of misinformation. "Hybrid" has gradually increasing shares for factual information, while "Community-Based" and "None" slowly decrease, though "None" has more overall shares. "Third-Party" experiences the most overall shares, and the sharpest increase in shares of factual information over time as well. Overall, strong fact-checking encourages users to amplify factual content, whereas misinformation rarely achieves viral spread across any setting.
Appendix L Emotional Salience of False vs. True News
To validate that our dataset reflects the known emotional drivers of rumor propagation, we conduct an auxiliary analysis comparing the emotional salience of real versus false news. Using GPT-4o as an annotator, we evaluated articles along established psychological dimensions of viral misinformation. We sampled 1,353 real news articles and paired them with 1,353 false articles from the NewsGuard dataset for a balanced 1:1 comparison.
Tab 5 reports average scores across six dimensions. False news consistently exhibits higher emotional intensity, fear, anger, and urgency, while being perceived as less credible.
Dimension Real False Difference Effect Size
Emotional Intensity Berger and Milkman (2012) 5.0 7.7 +2.7 +55%
Fear/Anxiety Vosoughi et al. (2018) 3.6 7.8 +4.2 +116%
Anger/Outrage Chuai and Zhao (2020) 3.1 7.6 +4.5 +145%
Urgency Song et al. (2023) 4.3 6.9 +2.6 +60%
Credibility Mang et al. (2024) 6.9 3.0 -3.8 -55%
Overall Rumor Potential Zhang et al. (2021) 4.0 8.5 +4.5 +110%
Table 5: Comparative emotional profile of real vs. false news articles (1–10 scale).
False news exhibits consistently higher levels of fear, anger, urgency, and emotional intensity, while being rated substantially lower in credibility. To complement this, Tab. 6 summarizes the proportion of articles that scored above 7/10 on each dimension. False news dominates across all measures, with particularly strong differences in anger and fear.
Metric Real News False News Ratio
High Emotional Intensity ($\geq$7/10) 335 (24.8%) 1,293 (95.6%) 3.9x
High Fear/Anxiety 236 (17.4%) 1,157 (85.5%) 4.9x
High Anger/Outrage 186 (13.7%) 1,170 (86.5%) 6.3x
High Rumor Potential 175 (12.9%) 1,310 (96.8%) 7.5x
Table 6: Distribution of high-emotion articles across real and false news.
These findings confirm that the NewsGuard dataset is well-calibrated to reflect the emotional salience typical of real-world misinformation. Specifically, false news demonstrates 7.5x higher viral potential, 145% greater anger/outrage, and 116% greater fear/anxiety than real news, consistent with prior work on rumor spread dynamics. This provides additional support for the realism of our simulation experiments.
Appendix M Prompts
M.1 Post Creation Prompt
@@
M.2 Feed Reaction Prompt